A box contains 79 red marbles, 76 white marbles, and 41 blue marbles. If a marble is randomly selected from the box, what is the probability that it is red or blue? Express your answer as a simplified fraction or a decimal rounded to four decimal places.
Answer #1

Total marbles = 79 + 76 + 41 = 196

Number of red or blue marbles = 79 + 41 = 120

So,

P(Red OR blue ) = N(Red OR blue) / Total marbles

= 120 / 196

= 30/49 ( = 0.6122 )